# Classification of Histopathology Slides with Persistence Homology Convolutions

### v1

@@ Abstract
Convolutional neural networks (CNNs) are a standard tool for computer vision tasks such as image classification. However, typical model architectures may result in the loss of topological information. In specific domains such as histopathology, topology is an important descriptor that can be used to distinguish between disease-indicating tissue by analyzing the shape characteristics of cells. Current literature suggests that reintroducing topological information using persistent homology can improve medical diagnostics; however, previous methods utilize global topological summaries which do not contain information about the locality of topological features. To address this gap, we present a novel method that generates local persistent homology-based data using a modified version of the convolution operator called Persistent Homology Convolutions. This method captures information about the locality and translation invariance of topological features. We perform a comparative study using various representations of histopathology slides and find that models trained with persistent homology convolutions outperform conventionally trained models and are less sensitive to hyperparameters. These results indicate that persistent homology convolutions extract meaningful geometric information from the histopathology slides.
keywords:
[…]
@@ 1 Introduction
With recent advances in machine learning, there has been a concurrent increase in the availability of large image data sets. Convolutional neural networks (CNNs) are a standard tool for tasks such as image classification, object detection, segmentation, and more. More recently, Vision Transformers (ViTs) have become a popular alternative to CNNs, achieving state of the art results in specific computer vision tasks [27, 28]. The utility of these models have been seen across numerous domains, especially in the field of computational medicine [25, 23, 24]. One such example is their use in both labeling and diagnosing diseases such as cancer [26].
@@
Both CNNs and ViTs have been trained to achieve a high degree of accuracy comparable to the performance of a trained pathologist [45, 12, 13, 47]. However, the CNN pooling operator may alter the relevant geometry within an image [2]. Similarly, the subdivision of images into patches for ViTs alters the geometric structure relative to each patch and may impact performance—for example, by subdividing a region with tumor growth. The geometric structure of cell tissue is an important characteristic in the field of histopathology: a branch of pathology that employs microscopy to examine tissue samples for disease-indicating abnormalities. In diseases like cancer, the cellular abnormalities are often geometric in nature with tissue samples showing varied cell and nucleus size, multinuclation, and a disorganization of tissue structure as illustrated in Figure 1.
@@
Figure 1: An example of differing tissue structure of between Non-Tumor (left), Necrotic Tumor (center), and Viable Tumor (right) samples in the Osteosarcoma image dataset from [4].
@@
It is natural to ask whether model performance can be improved by reincorporating a summary of the geometry. These data should reflect key information about the architecture of the tissue and shape of the cells. Geometric summaries can be created by utilizing tools from a branch of mathematics known as applied and computational topology; topology is the study of how geometric properties of a space change under continuous transformations. One focus of that field is the development of methods to perform shape analysis (e.g., determine the number of connected components, holes, cavities, etc) [5]. In recent decades, topology has emerged as an effective tool in data science and deep learning [37]. In particular, Persistent Homology (PH) can be used to represent geometric features of low-dimensional data and detect the topology of high-dimensional data sets. PH was discovered independently by Forsini and Landi [20], Robins [19], and Edelsbrunner et al [8, 10, 9]. One appeal of PH is that it provides a vectorizable representation of the shape of data; a perspective that is relevant to histopathology as demonstrated by Lawson et al [30]. In that specific study, the authors use the persistent homology of a sublevel filtration to quantify the cellular architecture of prostate cancer to accurately predict the Gleason score. Numerous studies have shown utility of PH in histopathology classification [38, 39, 40, 30].
@@
Qaiser et al. demonstrated that CNN model performance in image classification tasks can be improved by combining PH data with image data, as opposed to training a model with image data alone [3]. In their study, PH was computed “globally” in the sense that it was measured at the level of an entire image rather than for smaller image patches. We propose that computing PH “locally” — that is, separately in multiple smaller image patches — may result in better performance by enabling the retention of information related to the relative placement of topological features and reduce computational run time. We hypothesize that the local arrangement of topological features is an important characteristic for the differentiation tissue structure, and that some of this information is not detected by the models. To illustrate this, consider Figure 2. As we explain below, the one-dimensional persistent homology of each image will contain an interval for each cell summarizing its size and shape. As the two images contain about the same numbers of large and small cells, their global PH will be quite similar. However, their local structure is distinct.
@@
Figure 2: Images showing the cellular structure of nonwoody stems in a squash (left, Cucurbita sp) and in a castor oil plant (right, Ricinus communis). The tissues contain similar numbers of smaller and larger cells in noticeably different arrangements. Images were taken from [22].
@@
A simple approach to computing local persistent homology data is to subdivide the images into fixed patches as done in a ViT and to perform PH computations on separately these subimages before inputting them into other machine learning pipelines for training. Since the ViT’s model architecture includes a self-attention mechanism this method will accurately describe the local topological features in the image data and their relationships to one another [1]. However, if a region with tumor growth is at the boundary of a patch, this subdivision may result in the loss of important geometric information. Instead, we propose to compute local persistence over a family of possible overlapping patches, similar to how stride is used with the convolution operator in a CNN. That is, the Persistent Homology Convolution (PHC) of an $N\times N$ array localized to an $M\times M$ subwindows of $X$ can be expressed as
$[X\star_{PH}k]=\sum_{i}{\sum_{j}{k(c\cdot i,c\cdot j)}\mathbf{vec}(\mathrm{PH}_{p}(\mathcal{F}(T_{(c\cdot i,c\cdot j)}(X)))}$
@@
where $c$ is the stride length, $T_{(c\cdot i,c\cdot j)}$ is the restriction of $X$ to an $M\times M$ subimage whose bottom left coordinate is $(c\cdot i,c\cdot j)$, $\mathcal{F}$ is a function assigning a filtration to an $M\times M$ subimage, and $\mathbf{vec}$ is a vectorization operator. More detail regarding this equation is provided in Sections 2 and 3. In short, PHC provide a summary of the geometry in an image that captures information about its locality and translation invariance.
@@
To illustrate the effectiveness of PHCs we conduct a comparative study of the performance of CNNs trained using multiple different image data representations. The models are trained on an osteosarcoma dataset to distinguish between Non-Tumor, Non-Viable Tumor, and Viable Tumor classes [4, 45]. We perform over 10,000 experiments varying image representations and hyperparameters. We find that models trained on PHCs consistently outperform those trained using other data representations on all metrics. Our best performing model yielded $93.8\%$ accuracy in slide-based classification compared to $91.2\%$ accuracy in a previous study on the same data set applying conventionally trained CNNs with a similar architecture [^1] They did achieve $93.3\%$ after increasing the sample size by subdividing the images; we have not measured the effect of this on the accuracy of our models. The same study also tried multiple other conventional machine learning techniques in conjunction with shape analysis resulting in accuracies between 80.2% to 89.9%.  [45] Our results suggest that persistent homology convolutions reduce the complexity of image data while retaining relevant information about the geometry of histopathology slides. The main contributions to the paper are as follows:
@@ 2 Background
2.1 Persistent Homology
@@
To summarize the geometry of a cell arrangement in a slide, we apply a hybrid approach that measures the geometric properties of topological features using persistent homology (PH). PH encodes both qualitative (e.g. connectedness of tissue and number of cells) and quantitative (e.g. height, width, area of cells) information about this geometry. It is computed by associating to a data set a sequence of spaces (called a filtration) and measuring how the topology changes through the sequence. There are several different filtrations that can be associated to a data set. Most classically, one can associate to a subset $X$ of the plane a sequence of growing neighborhoods $X_{\epsilon}$ consisting of all points within distance $\epsilon$ of $X.$ The persistent homology of $X$ will then measure how components of $X$ merge and how holes in $X$ form and are filled in as $\epsilon$ increases. For point cloud data in low-dimensional Euclidean space, this can be computed in practice using the alpha complex [44]. Alternatively, the lower star filtration of a grayscale image $X$— denoted $\mathrm{St}_{\underline{\hskip 5.69054pt}}(X)$ —is the sequence of sets $X_{\rho}$ consisting of the union of all pixels whose grayscale value is less than or equal to $\rho.$ This filtration was applied to histopathological classification in [30].
We found that a third filtration yields better results for histopathology classification. Specifically, we compute what is known as the extended persistence of the height function $f$ (height PH) of the images to analyze the tissue architecture in the histopathology slides. This is done by scanning the images from bottom to top to observe when individual features are born and when they merge with other feature or die. That is, we associate to the image $X$ the filtration of level sets $f^{-1}(-\infty,a]=\{x\in X\colon f(x)\leq a\}.$ In the top left of Figure 3, this is shown by sweeping up a bar to reveal more and more of the image. Then, the process is continued by sweeping down the image in a different manner. Rather than concealing or revealing more of the image, we collapse the level set $f^{-1}[a,\infty)=\{x\in X:f(x)\geq a\}$ to a single point. In Figure 3, this is depicted by filling in the part of the image above the bar (which has the same topological effect).
Figure 3: Left: a simplified cellular structure. Right: The extended persistence diagram, $\mathrm{Ext}_{p}(f)$, for $p=0,1$ for the structure on the left. Ascending the cellular structure, the diagram measures Type 1 and 2 features. Upon descent, Type 3 feature are measured. Each point in the diagram is color coded to the corresponding feature.
We provide an explicit geometric description of the extended persistence diagram $\mathrm{Ext}_{p}(f)$ for two-dimensional images which does not require knowledge of terms from algebraic topology. The general definition of the extended persistent homology of a function is much more involved, and can be found in [34, 5, 21].
Definition 1.
An Extended Persistent Diagram, denoted $\mathrm{Ext}_{p}(f)$, is a multiset whose elements consist of intervals of the form $(b,d)$ where $f$ is the height function on the image and $p$ is the dimensionality of the features recorded. For each interval, $b$ represents the height that a feature is born and $d$ is when it dies. Every interval in the diagram is one of three types, the first two of dimension $p=0$ and the third of dimension $p=1$:
@@
    These intervals correspond to connected components of the level sets of $f$ which eventually merge with another component (one that was detected earlier). At height $b$, a new component of $f^{-1}(\infty,b]$ enters the frame. It merges with another component at height $d.$
@@
    These intervals correspond to connected components of the entire space $X.$ $b$ and $d$ are the maximum and minimum $y$-coordinates of the component, respectively.

### v3

# Classification of Histopathology Slides with Persistent Homology Convolutions

@@ Abstract
Convolutional neural networks (CNNs) are a standard tool for computer vision tasks such as image classification. However, typical model architectures may result in the loss of topological information. In specific domains such as histopathology, topology is an important descriptor that can be used to distinguish between disease-indicating tissue by analyzing the shape characteristics of cells. Current literature suggests that reintroducing topological information using persistent homology can improve medical diagnostics; however, previous methods utilize global topological summaries which do not contain information about the locality of topological features. To address this gap, we present a novel method that generates local persistent homology-based data using a modified version of the convolution operator called Persistent Homology Convolutions. This method captures information about the locality and translation equivariance of topological features. We perform a comparative study using various representations of histopathology slides and find that models trained with persistent homology convolutions outperform conventionally trained models and are less sensitive to hyperparameters. These results indicate that persistent homology convolutions extract meaningful geometric information from the histopathology slides.
@@ 1 Introduction
With recent advances in machine learning, there has been a concurrent increase in the availability of large image data sets. Convolutional neural networks (CNNs) are a standard tool for tasks such as image classification, object detection, segmentation, and more. More recently, Vision Transformers (ViTs) have become a popular alternative to CNNs, achieving state of the art results in specific computer vision tasks (Wu et al., 2021b; a). The utility of these models have been seen across numerous domains, especially in the field of computational medicine (Farooq et al., 2017; Bullock et al., 2019; Dozen et al., 2020). One such example is their use in both labeling and diagnosing diseases such as cancer (Jinnai et al., 2020).
@@
Both CNNs and ViTs have been trained to achieve a high degree of accuracy comparable to the performance of a trained pathologist (Arunachalam et al., 2019; Komura and Ishikawa, 2018; Melanthota et al., 2022; Mall et al., 2023). However, the CNN pooling operator may alter the relevant geometry within an image (Kepp et al., 2019). Similarly, the subdivision of images into patches for ViTs alters the geometric structure relative to each patch and may impact performance—for example, by subdividing a region with tumor growth. The geometric structure of cell tissue is an important characteristic in the field of histopathology: a branch of pathology that employs microscopy to examine tissue samples for disease-indicating abnormalities. In diseases like cancer, the cellular abnormalities are often geometric in nature with tissue samples showing varied cell and nucleus size, multinuclation, and a disorganization of tissue structure as illustrated in Figure 1.
@@
Figure 1: An example of differing tissue structure of between Non-Tumor (left), Necrotic Tumor (center), and Viable Tumor (right) samples in the Osteosarcoma image dataset from (Leavey et al., 2019; Arunachalam et al., 2019).
@@
It is natural to ask whether model performance can be improved by reincorporating a summary of the geometry. The data should reflect key information about the architecture of the tissue and shape of the cells. Geometric summaries can be created by utilizing tools from a branch of mathematics known as applied and computational topology; topology is the study of how geometric properties of a space change under continuous transformations. One focus of that field is the development of methods to perform shape analysis (e.g., determine the number of connected components, holes, cavities, etc) (Edelsbrunner and Harer, 2022). In recent decades, topology has emerged as an effective tool in data science and deep learning (Carlsson, 2009). In particular, Persistent Homology (PH) can be used to represent geometric features of low-dimensional data and detect the topology of high-dimensional data sets. PH was discovered independently by (Frosini et al., 1999), (Robins, 1999), and (Edelsbrunner et al., 2002; Zomorodian and Carlsson, 2004; Edelsbrunner et al., 2008). One appeal of PH is that it provides a vectorizable representation of the shape of data; a perspective that is relevant to histopathology as demonstrated by (Lawson et al., 2019). In that specific study, the authors use the persistent homology of a sublevel filtration to quantify the cellular architecture of prostate cancer to accurately predict the Gleason score. Numerous studies have shown utility of PH in histopathology classification (Abousamra et al., 2023; Chittajallu et al., 2018; Vipond et al., 2021; Lawson et al., 2019).
@@
(Qaiser et al., 2019) demonstrated that CNN model performance in image classification tasks can be improved by combining PH data with image data, as opposed to training a model with image data alone. In their study, PH was computed “globally” in the sense that it was measured at the level of an entire image rather than for smaller image patches. We propose that computing PH “locally” — that is, separately in multiple smaller image patches — may result in better performance by enabling the retention of information related to the relative placement of topological features and reduce computational run time. We hypothesize that the local arrangement of topological features is an important characteristic for the differentiation of tissue structure, and that some of this information is not detected by the models. To illustrate this, consider Figure 2. As we explain below, the one-dimensional persistent homology of each image will contain an interval for each cell summarizing its size and shape. As the two images contain about the same numbers of large and small cells, their global PH will be quite similar. However, their local structure is distinct.
@@
Figure 2: Images showing the cellular structure of nonwoody stems in a squash (left, Cucurbita sp) and in a castor oil plant (right, Ricinus communis). The tissues contain similar numbers of smaller and larger cells in noticeably different arrangements. Images were taken from (Berkshire Community College, 2024).
@@
A simple approach to computing local persistent homology data is to subdivide the images into fixed patches as done in a ViT and to perform PH computations on separately these subimages before inputting them into other machine learning pipelines for training. Since the ViT’s model architecture includes a self-attention mechanism this method will accurately describe the local topological features in the image data and their relationships to one another (Dosovitskiy et al., 2021). However, if a region with tumor growth is at the boundary of a patch, this subdivision may result in the loss of important geometric information. Instead, we propose to compute local persistence over a family of possible overlapping patches, similar to how stride is used with the convolution operator in a CNN. That is, the Persistent Homology Convolution (PHC) of an $N\times N$ array localized to an $M\times M$ subwindows of $X$ can be expressed as
$\mathbf{K}\star_{PH}X]=\sum_{i}{\sum_{j}{\mathbf{K}\star}\mathbf{vec}(\mathrm{PH}_{p}(\mathcal{F}(T_{(c\cdot i,c\cdot j)}(X)))}.$
@@
Here, $c$ is the stride length, $T_{(c\cdot i,c\cdot j)}$ is the restriction of $X$ to an $M\times M$ subimage whose bottom left coordinate is $(c\cdot i,c\cdot j)$, $\mathcal{F}$ is a function assigning a filtration to an $M\times M$ subimage, $\mathbf{K}$ is a kernel matrix, and $\mathbf{vec}$ is a vectorization operator. More detail regarding this equation is provided in Sections 2 and 3. In short, PHC provide a summary of the geometry in an image that captures information about its locality and translation equivariance.
@@
To illustrate the effectiveness of PHCs we conduct a comparative study of the performance of CNNs trained using multiple different image data representations. The models are trained on an osteosarcoma dataset to distinguish between Non-Tumor, Non-Viable Tumor, and Viable Tumor classes (Leavey et al., 2019; Arunachalam et al., 2019). We perform over 10,000 experiments varying image representations and hyperparameters. We find that models trained on PHCs consistently outperform those trained using other data representations on all metrics. Our best performing model yielded $93.8\%$ accuracy in slide-based classification compared to $91.2\%$ accuracy in a previous study on the same data set applying conventionally trained CNNs with a similar architecture [^1] They did achieve $93.3\%$ after increasing the sample size by subdividing the images; we have not measured the effect of this on the accuracy of our models.. The same study also tried multiple other conventional machine learning techniques in conjunction with shape analysis resulting in accuracies between 80.2% to 89.9%  (Arunachalam et al., 2019). Our results suggest that persistent homology convolutions reduce the complexity of image data while retaining relevant information about the geometry of histopathology slides. The main contributions to the paper are as follows:
@@ 2 Background
2.1 Persistent Homology of Histopathology Slides
@@
To summarize the geometry of a cell arrangement in a slide, we apply a hybrid approach that measures the geometric properties of topological features using persistent homology (PH). PH encodes both qualitative (e.g. connectedness of tissue and number of cells) and quantitative (e.g. height, width, area of cells, color) information about this geometry. It is computed by associating to a data set a sequence of spaces (called a filtration) and measuring how the topology changes through the sequence. There are several different filtrations that can be associated to a data set.
Figure 3: Top left: a subimage of a necrotic tumor slide. Bottom left: the threshold set $X$ from the lighter regions around the cells. Right: a simplified representation of $X$; $\mathrm{TPH}_{p}(X)$ is approximated by taking the alpha complex of $X.$ A visualization of the topological features that emerge is shown for the thickening filtration on the sample (left) and the topologically equivalent alpha complex (right). Type 1 features emerge when connected components merge as $\epsilon$ increases. Type 2 features are born when holes form and die in the union of balls centered at the finite point sample.
2.1.1 The Thickening Filtration and the Alpha Complex
Most classically, one can associate to a subset $X$ of the plane a sequence of growing neighborhoods $X_{\epsilon}$ consisting of all points within distance $\epsilon$ of $X.$ The persistent homology of $X$ will then measure how components of $X$ merge and how holes in $X$ form and are filled in as $\epsilon$ increases. In this context, persistent homology measures the birth and death of topological features. For point cloud data in low-dimensional Euclidean space, this can be computed in practice using the alpha complex (Edelsbrunner and Mücke, 1994).
To apply this to a grayscale image, the image is thresholded to yield a subset $X$ of the unit square $\left[0,1\right].$ $X$ consists of a union of pixels. In practice, the persistent homology of the alpha complex of the set of centers of these pixels yields a collection of intervals that is, in a well-defined sense (Cohen-Steiner et al., 2005), very close to the persistent homology of the thickening filtration $\left\{X_{\epsilon\geq 0}\right\}$ of $X.$ We provide an informal, geometric description of this persistent homology specialized to this setting.
Proposition 1.
The Persistent Homology Diagram of the thickening filtration of $X\subset\left[0,1\right]^{2}$ is a multiset whose elements consist of intervals of the form $(b,d)$. For each interval, the initial coordinate $b$ represents the value of $\epsilon$ at which a feature is born and the second coordinate $d$ is when it dies. There are two distinct types of features measured by $\mathrm{TPH}_{p}(X)$. The former is of dimension $p=0$ and the latter is of dimension $p=1$:
@@ 2.1.1 The Thickening Filtration and the Alpha Complex
    An interval $(0,d)\in\mathrm{TPH}_{p}(X)$ for each component of $X.$ The component merges with another in the thickened set $X_{d}.$ A single component of $X$ lasts throughout the filtration and has $d=\infty.$
@@
    An interval $(b,d)\in\mathrm{TPH}_{p}(X)$ for each bounded component of the complement that appears in the thickened sets $X_{\epsilon}.$ The bounded components of $\mathbb{R}^{2}\setminus X$ correspond to intervals with $b=0.$ An birth time $b$ of an interval $\left(b,d\right)$ with $b>0$ corresponds to the time when a “gulf” of $X$ closes to form a component of $\mathbb{R}^{2}\setminus X_{b}.$ The component disappears in the thickened set $X_{d}.$ Roughly speaking, the death time $d$ measures the largest enclosed ball that can fit inside the gulf or hole of $X$ corresponding to the feature.
In our study, $X$ is the subset of pixels of a histopathology slide whose grayscale values are lighter than a fixed threshold. For example, the lighter regions in the right most slide of Figure 1 correspond to the space between cells, isolated by surrounding tissue matter. As such, Type 1 features measure the density of cells and the uniformity of their arrangement. Type 2 features correspond to cells or subcellular structures that are darker than the neighboring matter (multinucleation); the birth and death values measure information about the size and shape of those features. As we describe above, we approximate $\mathrm{TPH}_{p}(X)$ using the alpha complex of a finite subset $\left\{x_{1},\ldots,x_{n}\right\}$ of $X$; the persistent homology of the alpha complex is the same as $\mathrm{TPH}_{p}(\left\{x_{1},\ldots,x_{n}\right\}).$ See Figure 3.
2.1.2 The Extended Lower Star Filtration
Alternatively, the lower star filtration of a grayscale image XX— denoted $\mathrm{St}_{\underline{\hskip 5.69054pt}}(X)$ —is the sequence of sets $X_{\rho}$ consisting of the union of all pixels whose grayscale value is less than or equal to $\rho$ (ranging from black pixels assigned 0 to white pixels assigned $1.$). That is, we associate to the image $X$ the filtration of level sets $X_{\rho}=\{x\in X\colon f(x)\leq\rho\}.$ In the bottom two rows of Figure 4, this is shown by progressively increasing the sublevels to observe when individual features (new components or holes) are born and when they merge with other feature or die. This filtration was applied to a classification task on prostate cancer histology slides in (Lawson et al., 2019). Additional features are captured by the extended persistence of this filtration, which scans “back down” through the superlevel sets $X^{\rho}$ the superlevel set $\{x\in X\colon f(x)\geq\rho\}$ of $f.$ More formally, the extended lower star filtration is defined by appending the quotient spaces $\left(\left[0,1\right]^{2}/X^{\rho}:1\geq\rho\geq 0\right)$ (in descending order) to the collection of sublevel sets, where $\left[0,1\right]^{2}/X^{\rho}$ is obtained from the square $\left[0,1\right]^{2}$ by collapsing the superlevel set $X^{\rho}$ to a point. See Figure 5.
We provide an informal, geometric description of the extended persistence diagram $\mathrm{XPH}_{p}(f).$ We specialize to the case where the domain of $f$ is the unit square; this avoids complications and allows for an account which does not not require knowledge of algebraic topology. The technical background is covered extensively elsewhere; see (Agarwal et al., 2004; Edelsbrunner and Harer, 2022; Cohen-Steiner et al., 2009; Turner et al., 2024) for a rigorous treatment of the extended persistent homology of a function and (Robins et al., 2011; Delgado-Friedrichs et al., 2014) for accounts specific to the sublevel set filtration of a grayscale image.
Consider an $N\times N$ grayscale function, and subdivide $\left[0,N\right]^{2}$ into $N^{2}$ unit squares. The image induces a step function $f:\left[0,N\right]^{2}\to\left[0,1\right]$ that assigns to each unit square the corresponding grayscale value. The Extended Persistent Homology of $f,$ denoted $\mathrm{XPH}_{p}(f)$, measures how the topology of sublevel sets $f^{-1}\left(\left[0,l\right]\right)$ and superlevel sets $f^{-1}\left(\left[s,1\right]\right)$ change with the parameters $l$ and $s.$ To give an elementary description, we replace $f$ with a smooth function $\tilde{f}$ whose persistent homology is infinitesimally close to that of $f.$ First, $f$ is not continuous so we linearly interpolate between different values of $f$ in an infinitesimal boundary of each unit square to yield a continuous function. Next, we round off the corners of the square $\left[0,1\right]^{2}$ to yield a domain $D$ with a smooth boundary. Finally, we infinitesimally perturb the function to obtain a smooth function $\tilde{f}:D^{2}\to\left[0,1\right]$ with no degenerate critical points (that is, so that the Hessian matrix is nonsingular at each interior critical point, and the single variable function obtained by restricting $\tilde{f}$ to the boundary has no inflection points). $\tilde{f}$ is called a Morse function. By applying an additional perturbation if necessary we may assume that $\tilde{f}$ takes distinct values at each critical point.
2.1.3 Ordinary Persistent Homology
We begin by describing the ordinary persistent homology of $\tilde{f}$ in terms of pairs of critical points. To aid visualization, it is useful to picture the sublevel sets $D_{l}$ as not living in the planar set $D$ but rather in the graph of $\tilde{f}$ in three dimensional space, as shown at the top right of Figure 4. As such, we sometimes refer to $\tilde{f}\left(x\right)$ as the “height” of $\tilde{f}$ at the point $x.$
First, the 0-dimensional persistent homology measures how components of the sublevel sets $D_{l}\coloneqq\tilde{f}^{-1}\left(\left[0,l\right]\right)$ appear and merge as $l$ increases from zero to one. A new component emerges at each local minimum of $\tilde{f}$, whether they are on the boundary or the interior. The component born at the absolute minimum $b_{0}$ of $\tilde{f}$ never disappears, so it corresponds to an an interval $\left(b_{0},\infty\right)$ in dimension $p=0.$ Other local minima are paired either with saddle points of $\tilde{f}$, as illustrated in the second row of Figure 4, or with local maxima on the boundary. This can be seen as follows. Suppose $x_{0}$ and $x_{1}$ are two local minima of $\tilde{f}$. Since $D_{1}$ is connected, there is an $0<l^{\prime}<1$ so that $x_{0}$ and $x_{1}$ are in different connected components of $D_{l}$ for $l<l^{\prime}$ but in the same component of $D_{l^{\prime}}.$ A path from $x_{0}$ to $x_{1}$ must pass through a point $x_{2}$ so that $\tilde{f}\left(x_{2}\right)=l^{\prime}.$ If $x_{2}$ is in the interior of $D$ then there must be directions in which $\tilde{f}$ is both increasing and decreasing at $x_{2}$ (if not, $x_{0}$ and $x_{1}$ would be connected in a lower level set) so $x_{2}$ must be a saddle point of $\tilde{f}$. Otherwise, if it is on the boundary it must be a local maximum.
Now, consider a local maximum $x$ of $\tilde{f}$ in the interior of its domain, such as the one illustrated in the bottom row of Figure 4. The sublevel set $D_{f\left(x\right)-\epsilon}$ contains a loop around $x$ which cannot be contracted to a point (otherwise, $\tilde{f}$ would not be a local maximum). This loop generates a one-dimensional homology class which “dies” at $d:=\tilde{f}\left(x\right).$ $D_{0}$ does not contain any cycles, so this one must be “born” at some point. If we replaced $\tilde{f}$ with $\tilde{f}$ the local minimum $x_{0}$ of $\tilde{f}$ would be paired with a saddle point (or boundary local maximum) of $\tilde{f}$ at some point $x_{1}.$ This is also a saddle point (or boundary local minumum) of $\tilde{f}$ and it turns out the the cycle which dies at $d$ first appeared at $b:=\tilde{f}\left(x_{1}\right),$ perhaps splitting from another loop.
Figure 4: Top left: histopathology slide of necrotic tumor cells. Top right: the graph of $\tilde{f}$ in the vicinity of a cell with multi-nucleation. $\tilde{f}$ possesses many minima, saddle, and maxima; pairs of neighboring critical points correspond to different types of persistent homology intervals. Each colored box on the density plot has a simplified representation of the geometric feature below that appears in the filtration. There is a single Type 1 feature (magenta box) corresponding to the absolute minimum of $\tilde{f}$. Type 2 features (black box) pair local minima with neighboring saddle points (or maxima on the boundary); they represent cellular or subcellular structures darker than neighboring pixels. The illustrated feature quantifies multi-nucleation. Type 3 features (blue box) pair local maxima with neighboring saddle points; these pairs capture measurements related to structures lighter than their neighbors. Here, a void between cells gives rise to a Type 3 feature.
Proposition 2.
[…]
  • Type 1.
[…]
  • Type 2.
[…]